OpenStudy (anonymous):

John rode his bike 2/3 miles to school, 3/5 miles to the mall, and 4/15 miles back home. What is the total distance he rode?

OpenStudy (anonymous):

\[\huge distance=\frac{ 2 }{ 3 }+\frac{ 3 }{ 5 }+\frac{ 4 }{ 15 }\]

OpenStudy (anonymous):

We need the denominators to all have the same number! In order to do that, we have to find a common multiple. We see that 15 is contains a factor of 3 and 5, so this is perfect! \[\huge distance=\frac{ 5 \cdot 2 }{ 5 \cdot3 }+\frac{ 3 \cdot 3 }{ 3 \cdot 5 }+\frac{ 4 }{ 15 }\]

OpenStudy (anonymous):

I'll let you do the rest :)